A wooden Japanese temple guardian statue of the Kamakura period (AD 1185 - 1334) showed a carbon-14 activity of 12.9 dpm/g-C when it was discovered in 1990. If the initial 14C activity was 15.0 dpm/g-C, what year was the statue made? ( t½ = 5730 yrs)


1
Expert's answer
2020-11-23T06:59:46-0500

Radioactive decay follows first order kinetics.Therefore,Decay constant (λ)=t1/2​0.693=57300.693​Given, [R0​]=100∴[R]=80and t=λ
2.303​log[R][R]0=(5730
0.693)
2.303​log80100=0.693
2.303×5730​×0.0969=1845 years.